OpenStudy (anonymous):

a = b a2 = ab a2 - b2 = ab-b2 (a-b)(a+b) = b(a-b) a+b = b b+b = b 2b = b 2 = 1 what is wrong here

OpenStudy (anonymous):

a+b = b b+b = b 2b = b 2 = 1

OpenStudy (anonymous):

the question is what is wrong in that proof ?

OpenStudy (anonymous):

a+b = b

OpenStudy (anonymous):

why ???

OpenStudy (anonymous):

if a = b, then both a + b = b?

OpenStudy (anonymous):

division by zero

OpenStudy (anonymous):

yes estudier got it !

OpenStudy (anonymous):

U have cancel (a-b) on both sides. Equivalent to division by zero since a=b

OpenStudy (anonymous):

Palmo4ka ur answers are smart